OpenStudy (anonymous):

$100 is deposited at the end of every week for five years in an account that pays 14%/a, compounded weekly.

OpenStudy (anonymous):

Find the future value of the annuity using the formula. Find the future value of the annuity using a spreadsheet. In your answer include the formulas that you typed. Find the furture value of the annuity using the TVM solver on a graphing calculator or on a website. In your answer, include what values you typed for each parameter.

OpenStudy (anonymous):

you need the annuity formula

OpenStudy (anonymous):

how do you find the compounding periods

OpenStudy (anonymous):

and what spreadsheet should i use?

OpenStudy (anonymous):

weekly means 52

OpenStudy (anonymous):

here is more descriptive formula $$\Large A = \frac{P~((1 + \frac r n )^{nt}-1)}{\frac r n }$$ P = regular payment into account r = annual interest rate n= # of times compounded a year t = # of years

OpenStudy (anonymous):

here r = 0.14 n = 52 t = 5

OpenStudy (anonymous):

okay thanks and what spreadsheet should i use?

OpenStudy (anonymous):

excel works :)

OpenStudy (wolf1728):

First you will need the compounded interest rate Rate = [(1 + (annual rate / 52) ^ 52] -1 Rate = [(1 + 14/52) ^52]-1 Rate = [(1.2692307692)^52]-1 Rate = 1.1500574252 -1 Rate = .1500574252 Rate = 15.00574252 %

OpenStudy (anonymous):

i meant template

OpenStudy (wolf1728):

The free to download Open Office Calc works just as well as Excel

OpenStudy (anonymous):

i dont know how to make it

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

wolf, what is that formula?

OpenStudy (wolf1728):

don't know how to make what?

OpenStudy (anonymous):

im supposed to use a spreadsheet but i dont know how to make one that would give my answers

OpenStudy (anonymous):

ill post an excel spreadsheet

OpenStudy (anonymous):

okay thanks

OpenStudy (wolf1728):

The formula is converting an annual interest rate into the rate after it is compounded. In that case it converts an annual rate into the rate if it were compounded weekly.

OpenStudy (anonymous):

so should i use what wolf said or jayzdd?

OpenStudy (wolf1728):

I'm trying to find a weekly annuity formula Here is one for monthly annuities:

OpenStudy (wolf1728):

Where n is the number of weeks and the interest rate is rate divided by 5200 for number of weeks and 100 tochange from per cent to decimal

OpenStudy (anonymous):

@jayzdd can i please get the excel spreadsheet?

OpenStudy (anonymous):

A=P ((1+rn)^nt −1)/rn using this formula i got 5.4870 is that right?

OpenStudy (anonymous):

yes i have it

OpenStudy (anonymous):

okay thank you!

OpenStudy (wolf1728):

GEEZ - I got kicked off the question and a LOT of calculations went with it ARRGGGGHHHH!!!!!!!!!!!!!!

OpenStudy (anonymous):

oh no :(

OpenStudy (anonymous):

I use a notepad file and copy paste any Latex , and text, periodically (no pun intended).

OpenStudy (anonymous):

whats happening? lol

OpenStudy (anonymous):

@wolf1728 what calculations, i am interested

OpenStudy (anonymous):

@jayzdd A=P ((1+rn)^nt −1)/rn using this formula i got 5.4870 is that right?

OpenStudy (wolf1728):

n = 260 weeks rate = 15.00574252% / 5200 = 0.0028857197 Total = 100* ((1+ 0.0028857197)^261 -1)/0.0028857197) -100 Total = 38861.2021295111 Total = 38,861.20 (rounded)

OpenStudy (anonymous):

thats incorrect it should be A=P ((1+r/n)^(n*t) −1)/(r/n)

OpenStudy (anonymous):

yeah that is the formula i used

OpenStudy (anonymous):

you need parentheses in the denominator

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

so whos answer should i use? @wolf1728 or @jayzdd

OpenStudy (anonymous):

i am using this annuities formula

OpenStudy (anonymous):

Note that, the equation in that pdf is valid when depositing at *end* of each period. This is stated in your problem 'end of every week'. If you deposit at the beginning of each week, you would have to use a similar but different formula.

OpenStudy (anonymous):

okay thanks

OpenStudy (wolf1728):

jayzdd Yes, my formula is slightly different - among other things it calculates 261 weeks versus 260 weeks in your formula.

OpenStudy (anonymous):

do you have documentation for your formula?

OpenStudy (wolf1728):

It is from my website http://www.1728.org/annuity3.htm I had to change it for this problem from months to weeks

OpenStudy (wolf1728):

I'm gonna get going 7:00 a.m. c ya

OpenStudy (anonymous):

I don't think thats right. I got a different answer maybe the discrepancy has to do with when you are depositing the amount, at the beginning or end of the period. your example is to the right

OpenStudy (anonymous):

@shair do you have any questions about the excel the last part of the assignment you need a TVM calculator. i think you can use excel as well

OpenStudy (anonymous):

no thanks for your help!!

OpenStudy (anonymous):

it says to include formulas...which formula?

OpenStudy (anonymous):

is it the one in fx bar?

OpenStudy (anonymous):

for the tvm questions this works: lise deposits $1000 at the end of each month for five years in an account that earns 12%/a, compounded quarterly. Using the TVM, calculate the accumulated value at the end of five years. Solution Before entering values we need to calculate the nominal monthly interest rate. We are given the quarterly interest rate in the question is: Quarterly Interest Rate===12%÷43%0.03 We can use this to determine the nominal monthly rate, i, is: (1+0.03)41.125508811.12550881−−−−−−−−−√121.0099016340.009901634i======(1+i)12(1+i)121+i1+ii0.99% Lastly, we need to determine the annual interest rate: Annual Interest Rate==0.99%×1211.88% Now we can use the TVM solver to solve using these steps: Open the TVM Calculator. Enter the following: Periods = 60 (there are 12 x 5 = 60 payment intervals) Rate = 11.88 (annual interest rate) PV = 0.00 PMT = –1 000 (this value is negative since the payment is an outflow of cash) FV = Leave this cell blank

OpenStudy (anonymous):

but when i divide 14% with 52 it gives me an awkward value

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

it says to include formula. does that mean the one in the fx bar?

OpenStudy (anonymous):

i have to go back here

OpenStudy (anonymous):

what TVM calculator are you using

OpenStudy (anonymous):

with the calculator i got 37328.69

OpenStudy (anonymous):

first test it with values that make sense

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

did you get this http://prntscr.com/9hnspl

OpenStudy (anonymous):

I entered PV =0 Rate = 14% PMT= -1000 Periods = 5*52 = 260 Weekly then click on FV

OpenStudy (anonymous):

no for the rate i pu 210.7 (11.60 * 14)

OpenStudy (anonymous):

and for periods i did 14*5 = 70

OpenStudy (anonymous):

are you doing a different problem?

OpenStudy (anonymous):

oops, should be -100 http://prntscr.com/9hnujw

OpenStudy (anonymous):

what is the question again? lets start over

OpenStudy (anonymous):

okay so, im supposed yo use a tvm calc for the same question

OpenStudy (anonymous):

$100 is deposited at the end of every week for five years in an account that pays 14%/a, compounded weekly

OpenStudy (anonymous):

I entered PV =0 Rate = 210.7% PMT= -1000 Periods = 5*14 = 260 Weekly

OpenStudy (anonymous):

Using this calculator: http://www.zenwealth.com/BusinessFinanceOnline/TVM/TVMCalculator.html PV =0 Rate = 14% PMT= -1000 Periods = 260

OpenStudy (anonymous):

rate is the annual interest rate

OpenStudy (anonymous):

sorry for periods i got 70

OpenStudy (anonymous):

Open the TVM Calculator. Enter the following: Periods = 60 (there are 12 x 5 = 60 payment intervals) Rate = 11.88 (annual interest rate) PV = 0.00 PMT = –1 000 (this value is negative since the payment is an outflow of cash) FV = Leave this cell blan

OpenStudy (anonymous):

i followed that^^^

OpenStudy (anonymous):

different TVM calculators are slightly different. so you have to figure out how it behaves

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so should i go with what u said?

OpenStudy (anonymous):

the calculator im following says to find the annual interest rate...

OpenStudy (anonymous):

so 14% isnt the annual interest

OpenStudy (anonymous):

14 is the annual interest rate

OpenStudy (anonymous):

usually the rate given in a TVM calculator means the annual interest rate

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Before entering values we need to calculate the nominal monthly interest rate. We are given the quarterly interest rate in the question is: Quarterly Interest Rate===12%÷43%0.03 We can use this to determine the nominal monthly rate, i, is: (1+0.03)41.125508811.12550881−−−−−−−−−√121.0099016340.009901634i======(1+i)12(1+i)121+i1+ii0.99% Lastly, we need to determine the annual interest rate: Annual Interest Rate==0.99%×1211.88% Now we can use the TVM solver to solve using these steps: Open the TVM Calculator. Enter the following: Periods = 60 (there are 12 x 5 = 60 payment intervals) Rate = 11.88 (annual interest rate) PV = 0.00 PMT = –1 000

OpenStudy (anonymous):

this is what i followed

OpenStudy (anonymous):

can u please review this

OpenStudy (anonymous):

Elise deposits $1000 at the end of each month for five years in an account that earns 12%/a, compounded quarterly. Using the TVM, calculate the accumulated value at the end of five years.

OpenStudy (anonymous):

for this questions

OpenStudy (anonymous):

Elise deposits $1000 at the end of each month for five years in an account that earns 12%/a, compounded quarterly. Using the TVM, calculate the accumulated value at the end of five years. PV = 0 Rate = 12% PMT= -1000 Periods = 5*4 = 20

OpenStudy (anonymous):

then why did they input Periods = 60 (there are 12 x 5 = 60 payment intervals) Rate = 11.88 (annual interest rate)

OpenStudy (anonymous):

i cant see what you're looking at

OpenStudy (anonymous):

who is 'they'

OpenStudy (anonymous):

in the answer it says the period is 60